### v2

@@ 1.2 Related Work
Online Metric Matching.
@@
A majority of literature for online metric matching focuses on the adversarial setting in balanced markets. For general metrics with adversarial servers and requests, [KP93], [KMV94], and [Rag16] respectively give $(2n-1)$-competitive deterministic algorithms, which is also shown to be tight. When randomization is allowed, an $O(\log^{3}n)$-competitive algorithm is proposed by [MNP06]; it is subsequently improved by [BBGN14] to $O(\log^{2}n)$, which remains the best upper bound. In the random-order model, the deterministic algorithm of [Rag16] admits a competitive ratio of $O(\log n)$, which is optimal even for randomized algorithms, and [GK19] show that the competitive ratio of the Greedy algorithm lies between $\Omega(n^{0.22})$ and $O(n)$.
@@
For the line metric with adversarial servers and requests, [KP93] show that the competitive ratio of Greedy is at least $12^{n}-$ 1, and a deterministic algorithm with a sublinear competitive ratio is known [ABN+19]. Furthermore, [NR17] present an $O(\log^{2}n)$-competitive deterministic algorithm, which is later shown to be $O(\log n)$-competitive [Rag18]. Moving to randomized algorithms, [GL12] propose an $O(\log n)$-competitive algorithm for the more general doubling metrics. For lower bounds, [FHK05] show that no deterministic algorithm can achieve a competitive ratio strictly better than $9.001$; it is then improved to $\Omega(\sqrt{\log n})$, which holds even for randomized algorithms [PS23].
Stochastic Online Metric Matching.
The stochastic online metric matching problem, where the requests are independently drawn from a known probability distribution, has attracted considerable attention in the past few years. We start with the references for balanced markets. For the Euclidean space $[0,1]^{d}$, where $d$ is a constant not scaling with $n$, [Kan22] proposes the Hierarchical Greedy algorithm with a constant competitive ratio when all servers and requests are uniformly distributed, and [CKKZ23] achieve nearly optimal regret guarantees when servers and requests are allowed to follow different distributions. [BFP23] show that Greedy is $O(1)$-competitive when servers and requests are uniformly distributed in $[0,1]$.
Motivated by the potential value of excess servers, [AALS22] initiate the study of unbalanced markets. When servers and requests are uniformly distributed in $[0,1]$, they show that the total matching cost of Greedy can be substantially improved by transitioning from balanced markets to unbalanced markets, and their results imply that Greedy is $O(\log^{3}n)$-competitive in unbalanced markets. [BFP23] subsequently pinpoint the competitive ratio of Greedy to be $O(1)$ under the same setting. Moreover, Hierarchical Greedy also admits a constant competitive ratio in unbalanced markets when servers and requests are uniformly distributed in $[0,1]^{d}$ for all $d\geq 1$ [Kan22].
@@
In contrast to the fruitful achievements when both servers and requests are stochastic, much less results are known for adversarial servers and stochastic requests. In balanced markets, [GGPW19] give an $O((\log\log\log n)^{2})$-competitive algorithm for general metrics and distributions, which further achieves a constant competitive ratio for tree metrics. Moreover, in balanced markets with requests uniformly distributed in $[0,1]$, [BFP23] show that Greedy is $\Theta(\log n)$-competitive and Hierarchical Greedy is $\Omega(n^{1/4})$-competitive.
Various papers study other aspects of stochastic online metric matching. [TTC94] show that Greedy admits a competitive ratio of $O(\sqrt{n})$ on the unit disk. In addition, [SNL+23] characterize the cost of Greedy in unbalanced markets with servers and requests uniformly distributed in $[0,1]$ using a partial differential equation approach.
@@
In recent years, due to the unparalleled impact of machine learning, the paradigm of designing online algorithms using externally provided predictions is of particular interest to overcome both information-theoretic and computational barriers. This paradigm suitably models the practical scenarios where future uncertainty is predictable according to prior knowledge, where predictions can be generated by machine learning models, human experts, historical data, heuristics, or even crowdsourcing. In particular, this line of literature usually assumes the provided predictions are untrusted with their error oblivious to algorithms, and algorithms are required to be consistent and robust. Besides, the performance of algorithms usually deteriorates smoothly as a function of the prediction error when the prediction accuracy decreases.
This paradigm has been incorporated into the study of various classical problems including online metric matching [ACE+23b], metrical task systems [ACE+23b, ACE+23a], online facility location [JLL+22, ACL+21], online bipartite matching [AGKK23, JM22], and caching [GPSS22, AGKP22, LV21], among others.
It is worth of mentioning that [ACE+23b] also study online metric matching with predictions. As the primary difference, we assume that the predictions are request locations and given at the beginning, whereas in their model, each arriving request is accompanied by a prediction of the behaviors of the optimal offline algorithm until the end of the current round, i.e., all servers matched by the optimal offline algorithm for the requests arrived so far. [^2] They allow the predictions to be inconsistent, i.e., each prediction may not be a superset of previous predictions. Moreover, their prediction error is defined against an optimal offline algorithm as the summation of prediction errors of all rounds, where the prediction error of each round is defined as the minimum cost of moving the predicted matched servers to the actual servers matched by the optimal offline algorithm until the end of this round. Hence, their results are in general incomparable with ours.
We note that concurrent and independent work by [CCM25] gives a consistent (but non-robust) algorithm for online metric matching with predictions that is very similar to ours. In particular, they assume that the prediction is given as a distribution over the metric space, and their algorithm draws $n$ independent samples from this distribution as predicted request locations, whereas we directly assume that $n$ points in the metric space are given as a prediction.
@@ 2 Model Setup
In our model, $n$ requests and $m$ servers are located in a metric space $(X,\delta)$, where $n\leq m$ and $\delta:X\times X\to{\mathbb{R}}_{\geq 0}$ is a cost function satisfying the triangle inequality. Denote the locations of servers as $S=\{s_{1},\ldots,s_{m}\}$ and the locations of requests as $R=\{r_{1},\ldots,r_{n}\}$. The servers are all available at the beginning with their locations known, and the requests arrive sequentially with their locations revealed only upon arrivals. When a new request arrives, we need to match it to an available server immediately and irrevocably, and then the matched server becomes unavailable. The cost of matching a server $s$ and a request $r$ is $\delta(s,r)$, and we aim to minimize the total matching cost.
Depending on the context, we will assume the locations of servers and requests are either adversarially chosen or independently drawn from a known distribution $\mathbb{D}$ over $X$. For all algorithm ${\mathcal{A}}$, set of servers $S$ with $|S|\geq n$, and set of requests $R$ with $|R|=n$, denote $\mathrm{cost}_{{\mathcal{A}}}(S,R)$ as the cost of ${\mathcal{A}}$ with the initial set of servers being $S$ and the set of requests being $R$; denote $\mathrm{cost}_{{\mathcal{A}}}(S,n)$ as the expected cost of ${\mathcal{A}}$ when the initial set of servers is $S$ and $n$ requests are randomly drawn from $\mathbb{D}$; denote $\mathrm{cost}_{{\mathcal{A}}}(n)$ as the expected cost of ${\mathcal{A}}$ when $|S|=|R|=n$ and all servers and requests are randomly drawn from $\mathbb{D}$.
Given a matching $M$ between two sets $S$ and $R$, for each matched element $s\in S$ (resp. $r\in R$), we use $M(s)$ (resp. $M(r)$) to denote the element in $R$ (resp. $S$) that $s$ (resp. $r$) is matched to. We define the smoothness property of distributions as follows, which requires distributions to admit an upper bounded density function.
Definition 1 ($\beta$-smoothness).
We say that a distribution $\mathbb{D}$ over $X$ is $\beta$-smooth if $\mathbb{D}$ admits a density function $\mu_{\mathbb{D}}$ such that $\mu_{\mathbb{D}}(x)\leq\beta$ for every $x\in X$.
@@ 2.1 Benchmark
We adopt the optimal offline cost, defined as the optimal matching cost without any future uncertainty, as our benchmark. For all set of servers $S$ and set of requests $R$ such that $|S|\geq|R|$, denote $\mathrm{OPT}(S,R)$ as the minimum cost of all matchings between $S$ and $R$ that match all requests. Given a set of servers $S$ with $|S|\geq n$ and a distribution $\mathbb{D}$ over $X$, the optimal offline cost on $S$ with $n$ requests randomly drawn from $\mathbb{D}$ is defined as $\mathrm{OPT}(S,n)=\mathbb{E}_{R\sim\mathbb{D}^{n}}\left[\mathrm{OPT}(S,R)\right]$, where $R\sim\mathbb{D}^{n}$ denotes that the requests in $R$ are independently drawn from $\mathbb{D}$. Furthermore, define $\mathrm{OPT}(n)=\mathbb{E}_{S,R\sim\mathbb{D}^{n}}\left[\mathrm{OPT}(S,R)\right]$.
We use competitive ratio and regret to measure the performance of online algorithms. When servers and requests are adversarial, for given metric space $(X,\delta)$, we say that the competitive ratio of an algorithm ${\mathcal{A}}$ is $\alpha$ if for all concerned instances [^3] We require $|S|=n$ when considering balanced markets and $n\leq|S|\leq O(n)$ for unbalanced markets. $(S,R)$, $\mathrm{cost}_{{\mathcal{A}}}(S,R)\leq\alpha\cdot\mathrm{OPT}(S,R)$. Similarly, when servers are adversarial and requests are stochastic, for given metric space $(X,\delta)$ and distribution $\mathbb{D}$, we say that the competitive ratio of an algorithm ${\mathcal{A}}$ is $\alpha$ if for all concerned instances $(S,n)$, $\mathrm{cost}_{{\mathcal{A}}}(S,n)\leq\alpha\cdot\mathrm{OPT}(S,n)$; moreover, define the regret of an algorithm ${\mathcal{A}}$ for a concerned instance $(S,n)$ as
$\mathrm{Regret}({\mathcal{A}};S,n)\triangleq\frac{1}{n}\left(\mathrm{cost}_{{\mathcal{A}}}(S,n)-\mathrm{OPT}(S,n)\right)\,.$
2.2 Optimal Offline Costs for Euclidean Space
A considerable portion of this paper focuses on Euclidean space. In this subsection, we assume the metric space to be $[0,1]^{d}$ with the Euclidean distance and bound the optimal offline costs for smooth distributions and general distributions. All the omitted proofs in this subsection can be found in Appendix A.
For an arbitrary distribution $\mathbb{D}$, the following upper bounds for $\mathrm{OPT}(n)$ are known from the optimal transport literature.
Lemma 1 (Theorem 3.11 in [HSM22]).
For all $d\geq 1$ and distribution $\mathbb{D}$ on $[0,1]^{d}$,
$\mathrm{OPT}(n)\leq\begin{cases}O(\sqrt{n})&d=1\\ O(\sqrt{n}\log n)&d=2\\ O(n^{1-\frac{1}{d}})&d\geq 3\end{cases}\,.$
When restricted to smooth distributions, we improve the upper bound for $d=2$ in Lemma 1 by generalizing the proof of the upper bound in the Ajtai-Komlós-Tusnády Matching Theorem [AKT84], which asserts that $\mathrm{OPT}(n)=\Theta(\sqrt{n\log n})$ for uniform distribution $\mathbb{D}$ and $d=2$.
@@ 2.2 Optimal Offline Costs for Euclidean Space
Assume $\mathbb{D}$ to be $\beta$-smooth for some constant $\beta>0$. For all $d\geq 1$,
$\mathrm{OPT}(n)\leq\begin{cases}O(\sqrt{n})&d=1\\ O(\sqrt{n\log n})&d=2\\ O(n^{1-\frac{1}{d}})&d\geq 3\end{cases}\,.$
Next, we establish lower bounds for smooth distributions in balanced markets.
@@
Assume $\mathbb{D}$ to be $\beta$-smooth for some constant $\beta>0$. For all $d\geq 1$,
$\mathrm{OPT}(n)\geq\begin{cases}\Omega(\sqrt{n})&d=1\\ \Omega(n^{1-\frac{1}{d}})&d\geq 2\end{cases}\,.$
The lower bound in Lemma 3 for $d=1$ follows from the quantile mapping and the well-known $\Omega(\sqrt{n})$ lower bound for the optimal offline cost with uniform distribution, and the lower bounds for $d\geq 2$ rely on nearest-neighbor-distance.
Finally, we present lower bounds for smooth distributions in unbalanced markets, which are proved by employing nearest-neighbor-distance as well.
Lemma 4.
Assume $\mathbb{D}$ to be $\beta$-smooth for some constant $\beta>0$. For all $d\geq 1$ and $S\subseteq[0,1]^{d}$ with $n\leq|S|\leq O(n)$, $\mathrm{OPT}(S,n)\geq\Omega(n^{1-\frac{1}{d}})$.
@@ 2.3 The Simulate-Optimize-Assign-Repeat Algorithm
The Simulate-Optimize-Assign-Repeat (SOAR) algorithm is proposed by [CKKZ23] under the setting in which both servers and requests are stochastic. The SOAR algorithm, whose guarantee is stated below, will serve as an important building block of our algorithms for stochastic online metric matching and be invoked in a black-box manner.
@@
Suppose that $m=n$ and $\mathbb{D}$ is a distribution on $X$, where the cost function $\delta$ is not required to satisfy the triangle inequality. When all servers and requests are independently drawn from $\mathbb{D}$, an algorithm exists with an expected cost of $\sum_{t=1}^{n}\frac{1}{t}\cdot\mathrm{OPT}(t)$.
3 A Generic Approach for Beyond the Worst-Case Analysis
In this section, we present a generic algorithmic approach for online metric matching to go beyond the worst-case analysis. Then, we will demonstrate how to apply this approach to leverage the provided distributional information or predicted request locations.

### v3

@@ 1.2 Related Work
Online Metric Matching in the Adversarial Setting.
@@
A majority of literature for online metric matching focuses on the adversarial setting and balanced markets. For general metrics, [KP93], [KMV94], and [Rag16] respectively give (2n1)(2n-1)-competitive deterministic algorithms, which is optimal. When randomization is allowed, an $O(\log^{3}n)$-competitive algorithm is proposed by [MNP06], which is subsequently improved to $O(\log^{2}n)$ [BBGN14]. In the random-order model, [Rag16] proposes an $O(\log n)$-competitive deterministic algorithm, which is optimal even for randomized algorithms, and [GK19] show that the competitive ratio of the greedy algorithm lies between $\Omega(n^{0.22})$ and $O(n)$.
@@
For the line metric, [KP93] show that the competitive ratio of the greedy algorithm is at least $2^{n}-1$, and a deterministic algorithm with a sublinear competitive ratio is known [ABN+19]. Furthermore, [NR17] present an $O(\log^{2}n)$-competitive deterministic algorithm, which is later shown to be $O(\log n)$-competitive [Rag18]. [GL12] propose an $O(\log n)$-competitive randomized algorithm for the more general doubling metrics. Regarding lower bounds, [FHK05] show that no deterministic algorithm can achieve a competitive ratio strictly better than $9.001$; it is then improved to $\Omega(\sqrt{\log n})$, which holds even for randomized algorithms [PS23].
Online Metric Matching in the Stochastic Setting.
We start by reviewing the literature for balanced markets. For the Euclidean metric $0,1]^{d}$, [Kan22] presents an $O(1)$-competitive algorithm for the uniform distribution, and [CKKZ23] give an algorithm achieving nearly optimal regret for general distributions. [BFP23] show that the greedy algorithm is $O(1)$-competitive for $[0,1]$ with the uniform distribution.
Motivated by the potential value of excess servers, [AALS22] initiate the study of unbalanced markets, for which they show that the greedy algorithm is $O(\log^{3}n)$-competitive for $[0,1]$ with the uniform distribution. This is subsequently improved to $O(1)$ [BFP23]. [Kan22] gives a $O(1)$-competitive algorithm for unbalanced markets and the Euclidean metric $0,1]^{d}$ with the uniform distribution.
Online Metric Matching in the Semi-Stochastic Setting.
@@
In contrast to the fruitful development in the stochastic setting, much less is known for the semi-stochastic setting, for which all existing results are for balanced markets. For general distributions, [GGPW19] give an $O((\log\log\log n)^{2})$-competitive algorithm for general metrics, which is $O(1)$-competitive for tree metrics. Moreover, for $[0,1]$ with the uniform distribution, [BFP23] show that the greedy algorithm is $\Theta(\log n)$-competitive.
@@
In recent years, due to the unparalleled impact of machine learning, the paradigm of designing online algorithms with access to externally provided predictions is of particular interest to overcome both information-theoretic and computational barriers. This paradigm suitably models the practical scenarios where future uncertainty is predictable according to prior knowledge, where predictions can be generated by machine learning models, human experts, historical data, heuristics, or even crowdsourcing. In particular, it is commonly assumed that the provided predictions are untrusted with their error oblivious to the algorithm, and the algorithm is desired to be consistent and robust. This paradigm has been incorporated into the study of various classical problems including online metric matching [ACE+23b], metrical task systems [ACE+23a, ACE+23b], online facility location [ACL+21, JLL+22], online bipartite matching [JM22, AGKK23], and caching [LV21, GPSS22, AGKP22], among others.
The paper most relevant to our work is [ACE+23b], which also studies online metric matching with predictions. As the primary difference, we assume that the prediction is a set of request locations and is given in advance, whereas they assume that each arriving request is accompanied by a prediction of the behaviors of the optimal offline algorithm until the end of the current period, i.e., all servers matched by the optimal offline algorithm to the requests arrived so far. [^2] They allow the predictions associated with different requests to be inconsistent, i.e., each prediction may not be a superset of previous predictions. Moreover, their prediction error is defined against an optimal offline algorithm as the summation of prediction errors of all periods, where the prediction error of each period is defined as the minimum cost of all perfect matchings between the predicted matched servers and the actual matched servers. Hence, their results are incomparable with ours.
@@ 2 Model Setup
In our model, $n$ requests $R=(r_{1},\ldots,r_{n})$ and $m$ servers $S=\{s_{1},\ldots,s_{m}\}$ are located in a metric space $(X,\delta)$ with $n\leq m$. The servers are all available at time $t=0$ with their locations known to the algorithm. At each time step $t\in[n]$, request $r_{t}$ arrives with its location revealed to the algorithm and needs to be matched immediately and irrevocably to a server $s$ that has not been matched, incurring a cost $\delta(s,r_{t})$.
Depending on the context, we will assume the locations of servers and requests are either adversarially chosen or independently drawn from a known distribution $\mathbb{D}$. In the adversarial setting, both $S$ and $R$ are adversarially chosen. In the semi-stochastic setting, $S$ is adversarial, while the requests in $R$ are independently drawn from $\mathbb{D}$}. In the stochastic setting, all servers in $S$ and requests in $R$ are independently drawn from $\mathbb{D}$}.
For all algorithm $\mathcal{A}$}, set of servers $S$ with $|S|\geq n$, and tuple of requests $R$ with $|R|=n$, denote $\mathrm{cost}({\mathcal{A}};S,R)$ as the expected cost of $\mathcal{A}$} under $S$ and $R$; denote $\mathrm{cost}_{{\mathbb{D}}}({\mathcal{A}};S,n)$ as the expected cost of $\mathcal{A}$} under $S$ and $n$ requests independently drawn from $\mathbb{D}$; denote $\mathrm{cost}_{{\mathbb{D}}}({\mathcal{A}};n)$ as the expected cost of $\mathcal{A}$} under $n$ servers and $n$ requests, all independently drawn from $\mathbb{D}$. Given a matching $M$ between two sets $S$ and $R$, for each matched element $s\in S$ (resp. $r\in R$), we use $M(s)$ (resp. $M(r)$) to denote the element in $R$ (resp. $S$) that $s$ (resp. $r$) is matched to.
We define the smoothness property for distributions as follows.
Definition 1 (Smoothness).
We say that a distribution $\mathbb{D}$ over $X$, which supports a uniform measure $\mathbb{U}$, is $\beta$-smooth for $\beta\geq 1$ if the corresponding measure $\mu_{{\mathbb{D}}}$ satisfies $\mu_{\mathbb{D}}(X^{\prime})\leq\beta\cdot\mathbb{U}(X^{\prime})$ for every measurable subset $X^{\prime}\subseteq X$.
@@ 2.1 Benchmark
We adopt the optimal offline cost, defined as the optimal matching cost without any future uncertainty, as our benchmark. For all set of servers $S$ and tuple of requests $R$ such that $|S|\geq|R|$, denote $\mathrm{OPT}(S,R)$ as the minimum cost of all matchings between $S$ and $R$ that match all requests. Given a set of servers $S$ with $|S|\geq n$ and a distribution $\mathbb{D}$ over $X$, define $\mathrm{OPT}_{\mathbb{D}}(S,n):=\mathbb{E}_{R\sim\mathbb{D}^{n}}\left[\mathrm{OPT}(S,R)\right]$ as the optimal offline cost on $S$ with $n$ requests in $R$ independently drawn from $\mathbb{D}$. Furthermore, define $\mathrm{OPT}_{\mathbb{D}}(n):=\mathbb{E}_{S\sim{\mathbb{D}}^{n},R\sim\mathbb{D}^{n}}\left[\mathrm{OPT}(S,R)\right]$.
We use competitive ratio and regret to measure the performance of online algorithms. In the adversarial setting, we say that an algorithm $\mathcal{A}$} is $\alpha$-competitive if for all $S$ and $R$, $\mathrm{cost}({\mathcal{A}};S,R)\leq\alpha\cdot\mathrm{OPT}(S,R)$. In the semi-stochastic setting, we say that an algorithm $\mathcal{A}$} is $\alpha$-competitive if for all $\mathbb{D}$}, $S$, and $n$, $\mathrm{cost}_{{\mathbb{D}}}({\mathcal{A}};S,n)\leq\alpha\cdot\mathrm{OPT}_{\mathbb{D}}(S,n)$; also, define the regret of $\mathcal{A}$} as
$\mathrm{Regret}_{\mathbb{D}}({\mathcal{A}};S,n):=\frac{1}{n}\left(\mathrm{cost}_{{\mathbb{D}}}({\mathcal{A}};S,n)-\mathrm{OPT}_{\mathbb{D}}(S,n)\right)\,.$
In the stochastic setting, we say that an algorithm $\mathcal{A}$} is $\alpha$-competitive if for all $\mathbb{D}$} and $n$, $\mathrm{cost}_{{\mathbb{D}}}({\mathcal{A}};n)\leq\alpha\cdot\mathrm{OPT}_{\mathbb{D}}(n)$.
2.2 Benchmark for Euclidean Metric
A considerable portion of this paper focuses on the Euclidean metric. In this subsection, we assume the metric space to be $0,1]^{d}$ with the Euclidean distance, and we bound the optimal offline costs for smooth distributions and general distributions.
For an arbitrary distribution $\mathbb{D}$, the following upper bounds for $\mathrm{OPT}_{\mathbb{D}}(n)$ are known from the optimal transport literature.
Lemma 1 ([Tal92, DY95, BL19]).
For every distribution $\mathbb{D}$ over $0,1]^{d}$,
$\mathrm{OPT}_{\mathbb{D}}(n)\leq\begin{cases}O(\sqrt{n})&d=1\\ O(\sqrt{n\log n})&d=2\\ O(n^{1-\frac{1}{d}})&d\geq 3\end{cases}\,.$
Next, we establish lower bounds for $\mathrm{OPT}_{\mathbb{D}}(n)$ with smooth distribution $\mathbb{D}$}. We defer the proof of Lemma 2 to Section A.1.
@@ 2.2 Benchmark for Euclidean Metric
Assume $\mathbb{D}$ to be a $\beta$-smooth distribution over $0,1]^{d}$ for $\beta\geq 1$. Then,
$\mathrm{OPT}_{\mathbb{D}}(n)\geq\begin{cases}\Omega(\beta^{-1}\sqrt{n})&d=1\\ \Omega(\beta^{-\frac{1}{d}}n^{1-\frac{1}{d}})&d\geq 2\end{cases}\,.$
The lower bound in Lemma 2 for $d=1$ follows from the quantile mapping and the well-known $\Omega(\sqrt{n})$ lower bound for the optimal offline cost with the uniform distribution, and the lower bound for $d\geq 2$ relies on nearest-neighbor-distance. For $d=2$, we fail to generalize the lower bound of $\Omega(\sqrt{n\log n})$ in the Ajtai-Komlós-Tusnády Matching Theorem, whose proof crucially exploits the structures of the uniform distribution. We remark that the optimal matching for $d=2$ exhibits the most intricate structure, and we refer to [Kan22, Tal22] for more thorough discussion. Hence, we leave the task of closing the gap for $d=2$ to future work.
Finally, we present lower bounds for unbalanced markets with smooth distributions, whose proof also employs nearest-neighbor-distance. We defer the proof of Lemma 3 to Section A.2.
@@
Assume $\mathbb{D}$ to be a $\beta$-smooth distribution over $0,1]^{d}$ for $\beta\geq 1$. Then, for all $d\geq 1$ and $S\subseteq[0,1]^{d}$ with $|S|=\kappa n$ and $\kappa\geq 1$, $\mathrm{OPT}_{\mathbb{D}}(S,n)\geq\Omega((\beta\kappa)^{-\frac{1}{d}}n^{1-\frac{1}{d}})$.
@@ 2.3 The Simulate-Optimize-Assign-Repeat Algorithm
The Simulate-Optimize-Assign-Repeat (SOAR) algorithm is proposed by [CKKZ23] in the stochastic setting and balanced markets. It will serve as an important building block of our algorithms in the semi-stochastic setting, and we state its guarantee as follows.
@@
Suppose that $m=n$ and $\mathbb{D}$ is an arbitrary distribution over $X$, where the cost function $\delta$ may not satisfy the triangle inequality. Then, when servers and requests are independently drawn from $\mathbb{D}$, there exists an algorithm with expected cost $\sum_{t=1}^{n}\frac{1}{t}\cdot\mathrm{OPT}_{\mathbb{D}}(t)$.
3 Generic Framework for Leveraging Predictions of Request Locations
In this section, we present a generic framework for online metric matching to design algorithms that leverage the provided prediction in the form of a set of request locations, which operates by enhancing an algorithm that does not use predictions. Then, we apply this framework to the semi-stochastic setting.